Return to question Required information [The following information applies to the questions displayed below.] Part 1 of 4 1.6eBook b. Compute Randolph Companys deferred income tax expense or benefit. Hint Ask Deferred income tax expense $ (40,800) PC. Compute Randolph Companys effective tax rate. (Round your answer to 2 decimal places. eBook Hint Effective tax rate Ask Pd. Provide a reconciliation of Randolph Companys effective tax rate with its hypothetical tax rate of 21 percent (Amounts to
Answer #1

Answer (a):

Current income tax expense $(105,000)

Working:

Given:

Taxable Income = $500,000

Tax rate = 21%

Current Income Tax expense = 500000 * -21% = ($105,000)

Answer (b):

Deferred income tax expense $ (25,200)

Working:

Favorable temporary difference = ($200,000)

Unfavorable temporary difference = $80,000

Net favorable temporary difference = ($120,000)

Deferred Tax expense (liability) = (120000) * 21% = $25,200

Answer (c):

Effective tax rate 16.28 %

Working:

Total income tax provision = 105000 + 25200 = $130,200

Pretax book income = $800,000

Effective Tax rate = 130200 / 800000 = 16.28%

Answer (d):

ETR reconciliation (in $) Income tax expense at 21% $ Tax benefit from permanent difference | $ Income tax provision $ 168,00

Working:

Income tax expense = 800000 * 21% = $168,000

Tax befit from permanent difference = 180000* 21% = $37,800

Tax benefit from permanent difference in % = 37800 / 800000 = 4.73%